# Fold catastrophe in breaking waves

### v1

@@ 4.3 Finite time blowup and breaking
The extension of the region of validity of the fast–slow model (26) beyond the breaking region, following the path $\Gamma_{*}$ of Eq. (36), to include higher-order effects on the surface-slope evolution, requires a multiple-scale perturbation approach to solve for the slow dynamics and for the dynamics of the Hessians of the pressure and velocity fields in order to evaluate the hotspot speed $C_{e}$ in Eq. (85) away from the marginal point $O$. Then, the fold catastrophe could be the cross-section of a cubic cusp catastrophe, or a higher-order swallowtail type (Thom1975; arnold1984catastrophe; zeeman1976catastrophe).
@@
Figure 4: Wave dynamics in the $(s,m)$ domain. Trajectories (black lines) of the four major successive waves in the simulated right-going wave group $C3N5A0514$, reported in BarthelemyJFM2018, are shown. These trajectories are labeled $0,1,2,3$ at their initial points (before focus) and are depicted with increasing thickness. The thickest (red) line corresponds to the breaking wave. The curves $\Gamma_{1}=\{(s,m):\tan\theta^{+}_{U}+m=0\rightarrow s=4(-m+m^{3})/(1-6m^{2}+m^{4})\}$ and $\Gamma_{2}=\{(s,m):m_{2}=0\rightarrow s=-m\}$, delimit the wave dynamics between the asymptotic slopes $\theta_{*}=\pm 22.5^{\circ}=\pm\arctan(\sqrt{2}-1)$. The curve $\Gamma_{*}={(s,m):s=(-m+5m^{3})/(1-6m^{2}+m^{4})}$ represents the approximate maximally non-breaking (marginal) wave. The breaking wave (red) exceeds the $\theta_{*}$ threshold, marking the onset of breaking inception.
@@
Figure 5: Eigen-structure of the weight matrices $G_{u}$ and $G_{p}$ associated with the Hessians of the fluid velocity $U$ and pressure $P$, respectively. Top panels: (Left) slope angle $\theta^{U}_{+}$ of the eigenvector of the velocity weight matrix $G_{u}$ corresponding to the positive eigenvalue $\lambda^{U}_{+}$ as a function of the surface slope $m$, for the four major successive waves of the simulated right-going wave group $C3N5A0514$ (BarthelemyJFM2018). (Right) Same, but for the slope angle $\theta^{P}_{+}$ of the eigenvector of the pressure weight matrix $G_{P}$ associated with the positive eigenvalue $\lambda^{P}_{+}$. The breaking wave is shown as the thickest dark-red line. Bottom panels: (Left) Ratio $\lambda_{+}^{U}/\lambda_{+}^{P}$ of the positive eigenvalues, and (right) ratio $\lambda_{-}^{U}/(-\lambda_{-}^{P})$ of the negative eigenvalues. The waves are labeled $0,1,2,3$ at the beginning (before focus) of their trajectories and are depicted with increasing thickness. The breaking wave is the thickest curve, colored distinctly from the non-breaking waves. The non-breaking waves are confined within the shaded region bounded by two limiting cases: (i) the condition $\tan\theta^{U}_{+}+m=0$, and (ii) vanishing momentum parameter $m_{2}=0$ (curves $\Gamma_{1}$, $\Gamma_{2}$ in Fig. 4). The dotted line represents the approximately maximally non-breaking (marginal) wave (curve $\Gamma^{*}$ in Fig. 4). The breaking wave surpasses the region bounded by the surface angle $\theta_{*}=\pm 22.5^{\circ}=\pm\arctan(\sqrt{2}-1)$ (dashed lines), marking the inception of breaking.
@@ 4.4 The point of no return: breaking inception
The breaking inception marks the point of no return, beyond which waves irreversibly break through a fold bifurcation, trespassing the the boundary $\partial\mathcal{F}_{0}$ at the onset of breaking (see Fig. 3). This boundary defines a skeleton in the domain $(m,A)$ that guides the wave dynamics near breaking. Interestingly, the stability of equilibria, expressed via $\partial_{m}F$, does not depend on $A$, which serves to define the fold boundary. Moreover, the point of breaking inception is robust and independent of how the waves are generated (DerakhtiJGR2020; BoettgerJFM2023; BoettgerPRF2024). In Appendix 12.6 we have derived the energetic hotspot speed that follows from Eq. (85) as
@@
with determinant $g_{p}=-(1+m^{2})(1+s^{2})/4$. These matrices encode the anisotropy of velocity and pressure fields and their effects on the wave surface dynamics. In particular, they weight the contribution of the two Hessians through their eigenvalues and eigendirections. Since the matrices are symmetric, their respective eigenvectors are also orthogonal. Their inverses are the matrix representations of pseudo-Riemannian, specifically hyperbolic metrics encoding anisotropy (see Appendix 12.6).
[…]
@@
Figure 4 shows the dynamics of the four major successive waves of the simulated right-going wave group $C3N5A0514$ (BarthelemyJFM2018). The non-breaking waves are confined within the area bounded by the two extreme paths $\Gamma_{1}$ and $\Gamma_{2}$, with the path $\Gamma_{*}$ representing the approximate maximally steep non-breaking (marginal) wave, close to wave $2$. The breaking wave $3$ trespasses the region delimited by the slope $\theta_{*}=\pm 22.5^{\circ}$ (dashed lines), which defines a breaking inception. Figure 5 summarizes the eigenstructure of the matrices $G_{u}$ and $G_{p}$ associated with the pressure $p$ and velocity $u$ Hessians tracked at the energetic hotspot. The point of breaking inception marks the maximal anisotropy between the velocity and pressure fields that the wave can sustain. This marks the dominance of the velocity Hessian at the breaking inception.
The role of the velocity and pressure Hessians in determining the breaking inception has a close analogue in turbulence. In incompressible flows, the pressure field is governed by a nonlocal Poisson equation, and its second derivatives enter the evolution equations for the velocity gradient tensor via the pressure Hessian (ohkitani1995pressureHessian; Chellard2006pressureHessian). The anisotropic part of the pressure Hessian prevents finite-time singularities by counteracting excessive alignment of strain and vorticity. Similarly, in the wave breaking problem, the weighted Hessians of pressure and velocity fields rule the approach to breaking.
@@
This trend is robust, as it consistently appears across other simulated breaking wave groups in BarthelemyJFM2018, indicating that the observed wave behavior and breaking inception is not case-specific but reflects a fundamental feature of the wave dynamics.
@@ 4.5 The $B_{+}$ and $B$ parameters
To characterize the inception of breaking at the energetic hotspot, BoettgerJFM2023 introduced the parameter
@@
At the energetic inception, we find $B_{+}^{*}\approx 0.6$ ($B\approx 0.55)$ with $s^{*}\approx 1$ and the surface slope $m^{*}=\sqrt{2}-1\approx 0.41$ both still significant away from the breaking onset (see Fig. 2). The momentum parameter $m_{2}^{*}=-1$ suggesting the vertical momentum transport is not dominant, while horizontal advection flows momentum toward the crest region. BoettgerJFM2023 also found that eventually the value of $B_{+}$ passes through $0.85$ at some time later prior to the breaking onset as the energetic hotspot gets closer to the crest. That is the same breaking inception threshold as observed by following the crest (BarthelemyJFM2018). Thus, the marginal point $O$ in Fig. 3, where the wave crest coincides with the energetic hotspot, suggests that the dynamics observed from these two points are intrinsically linked to the onset of breaking. In particular, the coincidence of crest and hotspot motion at $O$ marks the onset to breaking: the wave transitions from reversible focusing dynamics to irreversible breaking via a fold catastrophe (Thom1975; zeeman1976catastrophe; arnold1984catastrophe).
Estimating the critical threshold $B_{+}^{*}$ from Eq. (85) is a non-trivial task. It requires extending the region of validity of the fast-slow model (26) beyond the breaking region (see Fig. 3). Nevertheless, we can derive lower bounds as follows. At the breaking inception, the wave is still in the growing phase, and the hotspot elevation $h$ satisfies $\mathrm{d}h/\mathrm{d}\tau\geq 0$ (see Eq. (57) in Appendix 12.2). This yields the lower bound $B^{*}_{+}\geq m^{*}/(m^{*}-s^{*})\approx 0.3$.
Figure 6: (Left) Vertical Lagrangian acceleration $a_{L}/g$ (solid lines) and theoretical predictions from the slow-fast model (dashed lines) at the energetic hotspot as function of the fast time $\tau$, (center) ratio $\alpha=gh/K_{e}$ and (right) $\epsilon=1/U_{z}$ versus the surface slope $m$ for the four successive waves of the simulated right-going wave group $C3N5A0514$ (BarthelemyJFM2018). These trajectories are labeled $0,1,2,3$ at their initial stages (before focus) and are depicted with increasing thickness. The thickest (red) line corresponds to the breaking wave. At the breaking inception, the surface slope corresponds to the angle $\theta_{*}=-22.5^{\circ}$. For the limiting steady Stokes wave $\alpha=1$.
@@ 4.6 Vertical Lagrangian acceleration
where the term between square parentheses is $O(\epsilon)$ and balances $U_{z}=1/\epsilon$ to produce a finite Lagrangian acceleration at breaking. The left panel of Fig. 6 depicts the vertical Lagrangian acceleration $a_{L}$ inferred from simulations. Note that it exceeds the Stokes limit $g/2$, and in fair agreement with the predicted theoretical acceleration $a_{L}$ in Eq. (38), with the exception of underpredicting the breaking case. The simulations suggest that $a_{L}\rightarrow-g$ as the wave approaches the breaking onset. This critical threshold was originally proposed by phillips1966dynamics, and was later shown by Bridges_2009 to mark the transition from elliptic to hyperbolic character of the Euler equations. Higher-resolution simulations are needed to better resolve the wave field and surface to prove or disprove that $a_{L}+g\rightarrow 0$ as breaking is approached.
@@ 4.7 Maximum Crest Height
For the limiting steady Stokes wave, $\partial_{t}\phi|_{Q}=-2K_{e}$ and $\alpha=1$, while the vertical Lagrangian acceleration $a_{L}^{Stokes}=-UU_{z}=-P_{z}-g=-g/2$, with $P_{z}=-g/2$ being the vertical pressure gradient at the crest (Longuet_Higgins_AccelerationsinSteepGravityWaves; Shemer2013nhess). For nonstationary waves, however, $\partial_{t}\phi$ is expected to exceed the steady Stokes value in magnitude, reflecting the enhanced local accelerations of the fluid near steep, evolving crests.
The center panel of Fig. 6 shows the hysteresis loops of $\alpha$ versus the surface slope $m$ for four successive waves of the simulated right-going wave group $C3N5A0514$ (BarthelemyJFM2018). The ratio $\alpha$ is significantly greater than the Stokes limit for all cases and is bounded from below by $\alpha^{*}\approx 1.6$, indicating that nonstationarity plays a dominant role near breaking. This corresponds to the maximum surplus of kinetic energy $K_{e}$ over potential energy $gh$ that a wave can sustain, beyond which breaking becomes inevitable. This trend is observed in other simulated wave groups, suggesting upper bounds on the crest height of oceanic rogue waves (see, e.g., Fedele2025RogueWaves and references therein). Higher resolution numerical simulations are required to confirm these bounds and the full evolution of waves approaching the onset of breaking and beyond.
@@ 4.8 Finite-amplitude steady Stokes waves
A linear wave group can thus be regarded as induced by a perturbation of the linear Stokes wave at $L$ by spectral broadening. We note that DeconinckStokes2023 induced breaking by perturbing finite-amplitude nonlinear Stokes waves. This suggests that the segment between $L$ and $O$ could represent the locus of nonlinear Stokes waves, with the limiting Stokes wave located at $O$.
If the energetic hotspot speed $C_{e}$ is defined as the energy flux velocity of the finite-amplitude Stokes wave (PeregrineThomas1979; LonguetHiggins1975), then Fenton1985 fifth-order solution gives the finite-amplitude Stokes waves lie on LOL-O emanating from $L$. However, proving that the highest (limiting) Stokes wave is at the tip $O$ is not straightforward. Its group velocity tends to a finite value (LonguetHigginsFox1978part2), but evaluating the limiting value of the velocity gradient $U_{z}$ and $A$ requires refined numerical computations such as those performed by Creedon_Deconinck_Trichtchenko_2022 and DeconinckStokes2023. At present, therefore, the representation of finite-amplitude Stokes waves by the segment LOL-O, with the limiting Stokes wave at the cusp point $O$ remains a conjecture to prove or disprove.
@@ 4.9 Parity-reflection symmetry
Geometrically, the vertical axis $m=0$ serves as a parity-symmetry line of the fold boundary $\mathcal{F}_{0}$. When crossing this axis, the stability of equilibria flips: stable (unstable) fixed points become unstable (stable). Such a parity-reflection symmetry may be removed by applying symmetry reduction techniques described in (ChaosBook). For instance, for waves propagating to the right (left), the trajectory $m(t),A(t))$ can be restricted to the half-domain $m<0$ ($m>0$). Whenever it crosses the line $m=0$, the signs of $m$ is flipped, thereby folding the trajectory back into the $m<0$ ($m>0$) half-domain (see right panel of Fig. 3).
Figure 7: Dynamics in the plane $(m_{2},m)$ as observed from the inflection point. The points $I_{1}$ and $I_{2}$ denote the Stokes limiting slopes $m_{b}=\mp\tan(\pi/6)$ for right-going (-) and left-going (++) waves, respectively. The thick black lines indicate the loci of stable and unstable fixed points of the slope equation (42). Trajectories (blue lines) of the four tallest successive waves of the right-going simulated wave group $C3N5A0.514$ reported in BarthelemyJFM2018 are shown with increasing thickness, with the thickest (red) line corresponding to the breaking wave. See also Figure 1, which depicts the surface and energy profiles of wave $2$ and the location of its inflection point.

### v2

@@ 4.3 Finite time blowup and breaking
The extension of the region of validity of the fast–slow model (26) beyond the breaking region, following the path $\Gamma_{*}$ of Eq. (36), to include higher-order effects on the surface-slope evolution, requires a multiple-scale perturbation approach to solve for the slow dynamics and for the dynamics of the Hessians of the pressure and velocity fields in order to evaluate the hotspot speed $C_{e}$ in Eq. (85) away from the marginal point $O$. Then, the fold catastrophe could be the cross-section of a cubic cusp catastrophe, or a higher-order swallowtail type (Thom, 1975; Arnold, 1992; Zeeman, 1976).
@@
Figure 4: Wave dynamics in the $(s,m)$ domain. Trajectories (black lines) of the four major successive waves in the simulated right-going wave group $C3N5A0514$, reported in Barthelemy et al. (2018), are shown. These trajectories are labeled $0,1,2,3$ at their initial points (before focus) and are depicted with increasing thickness. The thickest (red) line corresponds to the breaking wave. The curves $\Gamma_{1}=\{(s,m):\tan\theta^{+}_{U}+m=0\rightarrow s=4(-m+m^{3})/(1-6m^{2}+m^{4})\}$ and $\Gamma_{2}=\{(s,m):m_{2}=0\rightarrow s=-m\}$, delimit the wave dynamics between the asymptotic slopes $\theta_{*}=\pm 22.5^{\circ}=\pm\arctan(\sqrt{2}-1)$. The curve $\Gamma_{*}={(s,m):s=(-m+5m^{3})/(1-6m^{2}+m^{4})}$ represents the approximate maximally non-breaking (marginal) wave. The breaking wave (red) exceeds the $\theta_{*}$ threshold, marking the onset of breaking inception.
@@
Figure 5: Eigen-structure of the weight matrices $G_{u}$ and $G_{p}$ associated with the Hessians of the fluid velocity $U$ and pressure $P$, respectively. Top panels: (Left) slope angle $\theta^{U}_{+}$ of the eigenvector of the velocity weight matrix $G_{u}$ corresponding to the positive eigenvalue $\lambda^{U}_{+}$ as a function of the surface slope $m$, for the four major successive waves of the simulated right-going wave group $C3N5A0514$ (Barthelemy et al., 2018). (Right) Same, but for the slope angle $\theta^{P}_{+}$ of the eigenvector of the pressure weight matrix $G_{P}$ associated with the positive eigenvalue $\lambda^{P}_{+}$. The breaking wave is shown as the thickest dark-red line. Bottom panels: (Left) Ratio $\lambda_{+}^{U}/\lambda_{+}^{P}$ of the positive eigenvalues, and (right) ratio $\lambda_{-}^{U}/(-\lambda_{-}^{P})$ of the negative eigenvalues. The waves are labeled $0,1,2,3$ at the beginning (before focus) of their trajectories and are depicted with increasing thickness. The breaking wave is the thickest curve, colored distinctly from the non-breaking waves. The non-breaking waves are confined within the shaded region bounded by two limiting cases: (i) the condition $\tan\theta^{U}_{+}+m=0$, and (ii) vanishing momentum parameter $m_{2}=0$ (curves $\Gamma_{1}$, $\Gamma_{2}$ in Fig. 4). The dotted line represents the approximately maximally non-breaking (marginal) wave (curve $\Gamma^{*}$ in Fig. 4). The breaking wave surpasses the region bounded by the surface angle $\theta_{*}=\pm 22.5^{\circ}=\pm\arctan(\sqrt{2}-1)$ (dashed lines), marking the inception of breaking.
@@ 4.4 The point of no return: breaking inception
The breaking inception marks the point of no return, beyond which waves irreversibly break through a fold bifurcation, trespassing the the boundary $\partial\mathcal{F}_{0}$ at the onset of breaking (see Fig. 3). This boundary defines a skeleton in the domain $(m,A)$ that guides the wave dynamics near breaking. Interestingly, the stability of equilibria, expressed via $\partial_{m}F$, does not depend on $A$, which serves to define the fold boundary. Moreover, the point of breaking inception is robust and independent of how the waves are generated (Derakhti et al., 2020; Boettger et al., 2023, 2024). In A.6 we have derived the energetic hotspot speed that follows from Eq. (85) as
@@
with determinant $g_{p}=-(1+m^{2})(1+s^{2})/4$. These matrices encode the anisotropy of velocity and pressure fields and their effects on the wave surface dynamics. In particular, they weight the contribution of the two Hessians through their eigenvalues and eigendirections. Since the matrices are symmetric, their respective eigenvectors are also orthogonal. Their inverses are the matrix representations of pseudo-Riemannian, specifically hyperbolic metrics encoding anisotropy (see A.6).
Near the breaking onset, where $\epsilon=1/U_{z}\ll 1$, from the scalings in Eqs. (102),(103), the energetic hotspot speed $C_{e}$ is nearly the same as the equivalent crest speed $C_{req}$, and the Hessian terms are negligible, that is
@@
Figure 4 shows the dynamics of the four major successive waves of the simulated right-going wave group $C3N5A0514$ (Barthelemy et al., 2018). The non-breaking waves are confined within the area bounded by the two extreme paths $\Gamma_{1}$ and $\Gamma_{2}$, with the path $\Gamma_{*}$ representing the approximate maximally steep non-breaking (marginal) wave, close to wave $2$. The breaking wave $3$ trespasses the region delimited by the slope $\theta_{*}=\pm 22.5^{\circ}$ (dashed lines), which defines a breaking inception. Figure 5 summarizes the eigenstructure of the matrices $G_{u}$ and $G_{p}$ associated with the pressure $p$ and velocity $u$ Hessians tracked at the energetic hotspot. The point of breaking inception marks the maximal anisotropy between the velocity and pressure fields that the wave can sustain. This marks the dominance of the velocity Hessian at the breaking inception.
The role of the velocity and pressure Hessians in determining the breaking inception has a close analogue in turbulence. In incompressible flows, the pressure field is governed by a nonlocal Poisson equation, and its second derivatives enter the evolution equations for the velocity gradient tensor via the pressure Hessian (Ohkitani and Kishiba, 1995; Chevillard and Meneveau, 2006). The anisotropic part of the pressure Hessian prevents finite-time singularities by counteracting excessive alignment of strain and vorticity. Similarly, in the wave breaking problem, the weighted Hessians of pressure and velocity fields rule the approach to breaking.
@@
This trend is robust, as it consistently appears across other simulated breaking wave groups in Barthelemy et al. (2018), indicating that the observed wave behavior and breaking inception is not case-specific but reflects a fundamental feature of the wave dynamics.
@@ 4.5 The $B_{+}$ and $B$ parameters
To characterize the inception of breaking at the energetic hotspot, Boettger et al. (2023) introduced the parameter
@@
At the energetic inception, we find $B_{+}^{*}\approx 0.6$ ($B\approx 0.55)$ with $s^{*}\approx 1$ and the surface slope $m^{*}=\sqrt{2}-1\approx 0.41$ both still significant away from the breaking onset (see Fig. 2). The momentum parameter $m_{2}^{*}=-1$ suggesting the vertical momentum transport is not dominant, while horizontal advection flows momentum toward the crest region. Boettger et al. (2023) also found that eventually the value of $B_{+}$ passes through $0.85$ at some time later prior to the breaking onset as the energetic hotspot gets closer to the crest. That is the same breaking inception threshold as observed by following the crest (Barthelemy et al., 2018). Thus, the marginal point $O$ in Fig. 3, where the wave crest coincides with the energetic hotspot, suggests that the dynamics observed from these two points are intrinsically linked to the onset of breaking. In particular, the coincidence of crest and hotspot motion at $O$ marks the onset to breaking: the wave transitions from reversible focusing dynamics to irreversible breaking via a fold catastrophe (Thom, 1975; Zeeman, 1976; Arnold, 1992).
Estimating the critical threshold $B_{+}^{*}$ from Eq. (85) is a non-trivial task. It requires extending the region of validity of the fast-slow model (26) beyond the breaking region (see Fig. 3). Nevertheless, we can derive lower bounds as follows. At the breaking inception, the wave is still in the growing phase, and the hotspot elevation $h$ satisfies $\mathrm{d}h/\mathrm{d}\tau\geq 0$ (see Eq. (57) in A.2). This yields the lower bound $B^{*}_{+}\geq m^{*}/(m^{*}-s^{*})\approx 0.3$.
Figure 6: (Left) Vertical Lagrangian acceleration $a_{L}/g$ (solid lines) and theoretical predictions from the slow-fast model (dashed lines) at the energetic hotspot as function of the fast time $\tau$, (center) ratio $\alpha=gh/K_{e}$ and (right) $\epsilon=1/U_{z}$ versus the surface slope $m$ for the four successive waves of the simulated right-going wave group $C3N5A0514$ (Barthelemy et al., 2018). These trajectories are labeled $0,1,2,3$ at their initial stages (before focus) and are depicted with increasing thickness. The thickest (red) line corresponds to the breaking wave. At the breaking inception, the surface slope corresponds to the angle $\theta_{*}=-22.5^{\circ}$. For the limiting steady Stokes wave $\alpha=1$.
@@ 4.6 Vertical Lagrangian acceleration
where the term between square parentheses is $O(\epsilon)$ and balances $U_{z}=1/\epsilon$ to produce a finite Lagrangian acceleration at breaking. The left panel of Fig. 6 depicts the vertical Lagrangian acceleration $a_{L}$ inferred from simulations. Note that it exceeds the Stokes limit $g/2$, and in fair agreement with the predicted theoretical acceleration $a_{L}$ in Eq. (38), with the exception of underpredicting the breaking case. The simulations suggest that $a_{L}\rightarrow-g$ as the wave approaches the breaking onset. This critical threshold was originally proposed by Phillips (1966), and was later shown by Bridges (2009) to mark the transition from elliptic to hyperbolic character of the Euler equations. Higher-resolution simulations are needed to better resolve the wave field and surface to prove or disprove that $a_{L}+g\rightarrow 0$ as breaking is approached.
@@ 4.7 Maximum Crest Height
For the limiting steady Stokes wave, $\partial_{t}\phi|_{Q}=-2K_{e}$ and $\alpha=1$, while the vertical Lagrangian acceleration $a_{L}^{Stokes}=-UU_{z}=-P_{z}-g=-g/2$, with $P_{z}=-g/2$ being the vertical pressure gradient at the crest (Longuet-Higgins, 1985; Shemer, 2013). For nonstationary waves, however, $\partial_{t}\phi$ is expected to exceed the steady Stokes value in magnitude, reflecting the enhanced local accelerations of the fluid near steep, evolving crests.
The center panel of Fig. 6 shows the hysteresis loops of $\alpha$ versus the surface slope $m$ for four successive waves of the simulated right-going wave group $C3N5A0514$ (Barthelemy et al., 2018). The ratio $\alpha$ is significantly greater than the Stokes limit for all cases and is bounded from below by $\alpha^{*}\approx 1.6$, indicating that nonstationarity plays a dominant role near breaking. This corresponds to the maximum surplus of kinetic energy $K_{e}$ over potential energy $gh$ that a wave can sustain, beyond which breaking becomes inevitable. This trend is observed in other simulated wave groups, suggesting upper bounds on the crest height of oceanic rogue waves (see, e.g., Knobler et al. (2025) and references therein). Higher resolution numerical simulations are required to confirm these bounds and the full evolution of waves approaching the onset of breaking and beyond.
@@ 4.8 Finite-amplitude steady Stokes waves
A linear wave group can thus be regarded as induced by a perturbation of the linear Stokes wave at $L$ by spectral broadening. We note that Deconinck et al. (2023) induced breaking by perturbing finite-amplitude nonlinear Stokes waves. This suggests that the segment between $L$ and $O$ could represent the locus of nonlinear Stokes waves, with the limiting Stokes wave located at $O$.
If the energetic hotspot speed $C_{e}$ is defined as the energy flux velocity of the finite-amplitude Stokes wave (Peregrine and Thomas, 1979; Longuet-Higgins, 1975), then Fenton (1985) fifth-order solution gives the finite-amplitude Stokes waves lie on LOL-O emanating from $L$. However, proving that the highest (limiting) Stokes wave is at the tip $O$ is not straightforward. Its group velocity tends to a finite value (Longuet-Higgins and Fox, 1978), but evaluating the limiting value of the velocity gradient $U_{z}$ and $A$ requires refined numerical computations such as those performed by Creedon et al. (2022) and Deconinck et al. (2023). At present, therefore, the representation of finite-amplitude Stokes waves by the segment LOL-O, with the limiting Stokes wave at the marginal point $O$ remains a conjecture to prove or disprove.
@@ 4.9 Parity-reflection symmetry
Geometrically, the vertical axis $m=0$ serves as a parity-symmetry line of the fold boundary $\mathcal{F}_{0}$. When crossing this axis, the stability of equilibria flips: stable (unstable) fixed points become unstable (stable). Such a parity-reflection symmetry may be removed by applying symmetry reduction techniques described in (Schneider et al., 2004). For instance, for waves propagating to the right (left), the trajectory $m(t),A(t))$ can be restricted to the half-domain $m<0$ ($m>0$). Whenever it crosses the line $m=0$, the signs of $m$ is flipped, thereby folding the trajectory back into the $m<0$ ($m>0$) half-domain (see right panel of Fig. 3).
Figure 7: Dynamics in the plane $(m_{2},m)$ as observed from the inflection point. The points $I_{1}$ and $I_{2}$ denote the Stokes limiting slopes $m_{b}=\mp\tan(\pi/6)$ for right-going (-) and left-going (++) waves, respectively. The thick black lines indicate the loci of stable and unstable fixed points of the slope equation (42). Trajectories (blue lines) of the four tallest successive waves of the right-going simulated wave group $C3N5A0.514$ reported in Barthelemy et al. (2018) are shown with increasing thickness, with the thickest (red) line corresponding to the breaking wave. See also Figure 1, which depicts the surface and energy profiles of wave $2$ and the location of its inflection point.